given market demand Q=100-P if the market supply function for the smaller firms is given by S=0.5P. And the cost function of the dominant firms is TC=10+40QD, where QD= output of the dominant firm. a) Find the market price and output of the dominant firm at equilibrium. b) Find the output level to be supplied by the smaller firms.


1
Expert's answer
2021-03-31T11:47:38-0400

Solution:

a.). Market price and output of the dominant firm:

At equilibrium: market demand = market supply

100 – P = 0.5P

100 = 0.5P + P

100 = 1.5P

P = "\frac{100}{1.5} = 67"


P = 67

Equate for Q:

Q = 100 – P

Q = 100 – 67

Q = 33

Market price = 67

Market quantity for the dominant firm = 33units

 

b.). Output level to be supplied by the smaller firms:

Smaller firms will supply output at the equilibrium price

Equilibrium price = 67

Smaller firms supply function: S = 0.5P

 = 0.5(67) = 33 units

Smaller firms will supply an output level of = 33 units